OpenStudy (princessaurora):

** WILL FAN & MEDAL ** Look at the system of equations: y = - x^2 + 1 and y = x^2. At which approximate points are the two equations equal?

OpenStudy (anonymous):

The y's are equal, right? That means you can set this up as x^2=-x^2 + 1. What should you do now?

OpenStudy (princessaurora):

Is it (-0.7,0.5)?

OpenStudy (anonymous):

Well, let's work it out. x^2=-x^2 + 1 2x^2=1 x^2=1/2 x≈0.7, -0.7 You can find the y-value by putting the solutions into y=x^2. .7^2≈.5 So they'll intersect at (0.7, 0.5) and (-0.7, 0.5)

OpenStudy (princessaurora):

ok i see but at what points are they equal?

OpenStudy (princessaurora):

both of those are answers

OpenStudy (princessaurora):

not one answer

OpenStudy (princessaurora):

on my worksheet

OpenStudy (anonymous):

can we see the multiple choice answers?

OpenStudy (anonymous):

Because there are two intersections. The question is probably asking for just one intersection, so yes, (-0.7, 0.5) would be the answer they're looking for. The intersections are where the equations are equal.

OpenStudy (princessaurora):

ok i see thanks !

OpenStudy (anonymous):

No problem!